OpenStudy (nosnip):

how to factorise and solve -x^2 - 8

OpenStudy (anonymous):

8 isn't a perfect square, so that can't be factored without using radicals. As for solving it, is that equal to something?

OpenStudy (nosnip):

that's what I thought but I've got to draw it as a graph

OpenStudy (nosnip):

there's y= (x-2)(x+2)(x+2) and y= -x^2 -8

OpenStudy (nosnip):

I've only got to sketch it but I need the x intercept I think

OpenStudy (anonymous):

so you need to sketch both of those and find the solution/intersection point?

OpenStudy (nosnip):

Yep

OpenStudy (anonymous):

ok. for the first one you can use the x-intercepts. The second one, doesn't have x-intercepts. You can draw it using transformations. Reflect the parabola over the x-axis and shift it down 8.

OpenStudy (nosnip):

Ooh I see yes that's clever!

OpenStudy (anonymous):

:) You can solve this algebraically as well \((x-2)(x+2)(x+2)=-x^2-8\) \((x^2-4)(x+2)=-x^2-8\) \(x^3+2x^2-4x-8=-x^2-8\) \(x^3+3x^2-4x=0\)

OpenStudy (nosnip):

Okay thankyo uvery much

OpenStudy (anonymous):

you're welcome